How many gallons of pure alcohol should be mixed with 20 gallons of a 15% alcohol solution to obtain a mixture that is 25% alcohol
Answer by mananth(16949)

How many gallons of pure alcohol should be mixed with 20 gallons of a 15% alcohol solution to obtain a mixture that is 25% alcohol
15% 20 gal
100% x
Mixture 25% x+20
..
0.15*20+x=0.25(x+20)
3+x=0.25x+5
add -0.25x to both sides
3+x-0.75x=0.25x-0.25x+5
0.25x+3=5
add -3 to both sides
0.25x +3-3=5-3
0.25x=2
divie by 0.25
x=2/0.25
x=8 gallons of pure alcohol